OpenStudy (anonymous):

LAST ONE FOR THE NIGHT! Use a graph to solve the equation on the interval [−2π, 2π]. tan x = - sqrt(3)

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@No.name

OpenStudy (anonymous):

@dan815

hero (hero):

Take the inverse tan of both sides \(x = \tan^{-1}(-\sqrt{3})\)

hero (hero):

You can do that to find one of the values. Afterwards add or subtract the period of tan to find the rest of the values.

OpenStudy (anonymous):

my calculator says error

hero (hero):

According to mine, there exists at least one value.

hero (hero):

Make sure you are in radian mode and make sure you are entering the values correctly.

OpenStudy (anonymous):

i am in radian mode. typing in \[\tan^-1(-\sqrt{3})\]

hero (hero):

You have to type \(\tan^{-1}(-\sqrt{3})\) precisely with no mistakes.

OpenStudy (anonymous):

that is exactly what I am typing in. let me try my online calculator

hero (hero):

What calc do you own?

OpenStudy (anonymous):

TI 83

OpenStudy (anonymous):

is it -pi/3?

hero (hero):

Put it this way, I have three different TI-calculators and I used all three of them and they all give me the same result. Yes -pi/3 is one of the values.

OpenStudy (anonymous):

is that our first interval? or our base?

hero (hero):

That's the starting point. Now add and subtract the appropriate periods of tan to find the rest of the values.

OpenStudy (anonymous):

subtract -2pi and add 2pi to -pi/3?

hero (hero):

What is the period of the tan function?

hero (hero):

Please answer the question first.

OpenStudy (anonymous):

isn't the period the 3 or sqrt(3)

hero (hero):

No, the period of the tangent function is \(\pi\). So keep adding and subtracting \(\pi\) from \(-\frac{\pi}{3}\) until you have all the values \(-2\pi\) < x < \(2\pi\)

OpenStudy (anonymous):

subtracting pi from -pi/3?

OpenStudy (anonymous):

subtract 5/3pi till -2pi

hero (hero):

Yes, that's right. You have to go in both directions because you have to find every value up to \(-2\pi\)

hero (hero):

You have to keep subtracting \(\pi\)

OpenStudy (anonymous):

add 7/3pi

OpenStudy (anonymous):

interval = -5/3pi, 7/3pi ???

hero (hero):

You're confusing yourself. I can feel it

hero (hero):

Keep adding and subtracting \(\pi\) only

OpenStudy (anonymous):

Right, these are my equations: -pi/3+pi(x)=2pi -pi/3-pi(x)=-2pi

hero (hero):

Once you find a positive value, add \(\pi\) to find the next value. Once you find a negative value, subtract \(\pi\) again to find the next value. You're supposed to find the values between -2\(\pi\) and 2\(\pi\) not set an equation equal to them.

OpenStudy (anonymous):

Im confused now. I am not sure what you mean. I thought you wanted to see how many pies we subtract from 5/3pi till you get -2pi. Then see how many pies we add to 5/3pi till you get 2pi.

OpenStudy (anonymous):

I would be incredibly grateful if you could give me the answers and then work from there?

hero (hero):

We found \(x = -\dfrac{\pi}{3}\). So we keep adding or subtracting \(\pi\) from that until we get to \(\pm 2\pi\).

OpenStudy (anonymous):

it does say "use a calculator" right?

OpenStudy (anonymous):

it says use a graph

OpenStudy (anonymous):

@Hero which is 7/3pi to get to 2pi

OpenStudy (anonymous):

@satellite73 that does not give us intervals

OpenStudy (anonymous):

so i am still doing this wrong?

OpenStudy (sweetburger):

well Hero cleared that up for me ty

hero (hero):

I meant to post this for a graph: https://www.desmos.com/calculator/dzd5suoweu

hero (hero):

The solutions occur where the red and green intersect if you know how to read the graph.